OpenStudy (anonymous):

Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April? A. 27 B. 35 C. 95 D. 103

OpenStudy (texaschic101):

j = a + 8 j + a = 198 now sub a + 8 in for j in the 2nd equation a + 8 + a = 198 2a + 8 = 198 2a = 198 - 8 2a = 190 a = 190/2 a = 95 now sub 95 in for a in either equation to find j j = a + 8 j = 95 + 8 j = 103 April is 95

OpenStudy (anonymous):

thank you so so so so much :)